OpenStudy (anonymous):

In a circle a 90 degree sector has area of 36 pi ft^2 what is the radius of the circle? Answers are either 18, 12, 6 or 16.

OpenStudy (anonymous):

area of 90 degree sector has 1/4 pieR^2=36 pi ft^2 1/4r^2=36 r^2=36*4 r=18

OpenStudy (phi):